# Kolmogorov-Arnold Attention: Is Learnable Attention Better For Vision Transformers?

### v1

@@ Appendix B Addendum to Empirical Analysis
B.1 Implementation Details
We implement the framework in Python using PyTorch (Paszke et al., 2019) and use the training strategy of DEiT (Touvron et al., 2021). We train all models for $100$ epochs with the ADAM optimizer (Kingma & Ba, 2015) with a base learning rate of $3.125\times 10^{-5}$ and batch-size $32$, except for the experiments on ImageNet-1K that use a learning rate of $1.25\times 10^{-4}$ with batch-size $128$. All experiments use a warm-up of $5$ epochs and a cosine scheduler with a weight decay of $5\times 10^{-2}$. The experiments were performed on two 80 GB NVIDIA H100 GPUs. All the hyper-parameter settings are given in Table 2.
Table 2: Hyper-parameter settings for all the experiments conducted in this work.
@@ B.1 Implementation Details
B.2 Computation Time and Throughput
@@
The overall computation for Fourier KArAt variants is higher than their conventional softmax MHSA, and we have delineated the same in Figure 10. Primarily, the Fourier KArAt variants have a longer training time. We show the training time comparison between the vanilla MHSA and its Fourier KArAt versions for 100 epochs on CIFAR-10, CIFAR-100, and Imagenet-1K datasets for all the models (ViT-Tiny, ViT-Small, ViT-Base) in Figures 101010, respectively. We only compare the best-performing Fourier KArAt models ($G_{1}B$ and $G_{1}U$ for ViT-Base, and $G_{3}B$ and $G_{3}U$ for ViT-Tiny & ViT-Small) with their softmax MHSA counterparts. We also observe that universal mode $G_{n}U$ training times are consistently slightly less than the blockwise modes $G_{n}B$.
@@
During the training, we monitored the GPU memory requirements, and as expected, Fourier KArAt variants utilize significantly more memory than vanilla MHSA. In particular, the GPU memory requirements scale by 2.5×2.5$\times$2.5 ×, 4.25×4.25$\times$4.25 ×, and 6×6$\times$6 × in case of ViT-Tiny, ViT-Small, and ViT-Base, respectively, compared to the vanilla softmax MHSA.
@@
We also compare the throughput during inference in Figure 10 and see slightly faster inference in universal mode than blockwise except for ViT-Base. While there is a massive training time discrepancy between vanilla ViTs and Fourier KArAt ViTs, the inference speeds for Fourier KArAt variants are comparable to their vanilla counterparts. Although there is a minor difference in throughput between universal mode and blockwise mode during inference, theoretically, both variants for any model with the same grid size should involve the same quantity of FLOPs.
@@
Figure 10: A detailed comparison of computing requirements. We compare the training times for 100 epochs with the hyperparameter settings given in Table 2 for all the datasets CIFAR10, CIFAR100, and ImageNet1K. We also compare the throughputs of different models on ImageNet1k; the throughput results will be similar for other datasets as the input size is $224\times 224$.
B.3 Ablation with the Hidden Dimension, $r$
While avoiding the computational overhead for computing $\Phi^{i,j}\in\mathbb{R}^{N\times N}$, we make use of the low-rank structure that the attention heads show (see Figure 3) by comparing different values of $r$. Particularly, we consider the values, $r=8,12,24$, on the Fourier KArAt variant of ViT-Base model, $\widehat{\Phi}^{i,j}\in\mathbb{R}^{r\times N}$; see Table 5 for results on CIFAR-10. In this ablation, we observe that changing the hidden dimension has a minimal impact on the model’s performance. This can be explained by the sudden drop in the singular values, as shown in Figure 7. As long as $r$ remains greater than the sudden drop index, the model should not be impacted by the changing of $r$ except for changes in computational requirement; a higher $r$ would incur a higher compute time as the size of the operator $\widehat{\Phi}$ scales with $r$.
     Model      Acc.@1
     CIFAR-10      CIFAR-100
     ViT-Base      83.45      58.07
     + $G_{1}B$      81.81      55.92
     + $G_{1}U$      80.75      57.36
     + $G_{3}B$      80.09      56.01
     + $G_{3}U$      81.00      57.15
     + $G_{5}B$      79.80      54.83
     + $G_{5}U$      81.17      56.38
     + $G_{11}B$      50.47      42.02
     + $G_{11}U$      40.74      39.85
     ViT-Small      81.08      53.47
     + $G_{1}B$      79.00      53.07
     + $G_{1}U$      66.18      53.86
     + $G_{3}B$      79.78      54.11
     + $G_{3}U$      79.52      53.86
     + $G_{5}B$      78.64      53.42
     + $G_{5}U$      78.75      54.21
     + $G_{11}B$      77.39      52.62
     + $G_{11}U$      78.57      53.35
     ViT-Tiny      72.76      43.53
     + $G_{1}B$      75.75      45.77
     + $G_{1}U$      74.94      46.00
     + $G_{3}B$      76.69      46.29
     + $G_{3}U$      75.56      46.75
     + $G_{5}B$      75.85      –
     + $G_{5}U$      74.71      –
     + $G_{7}B$      75.11      –
     + $G_{7}U$      74.45      –
     + $G_{9}B$      74.85      –
     + $G_{9}U$      73.97      –
     + $G_{11}B$      74.52      –
     + $G_{11}U$      73.58      –
Model $r$ Acc.@1 on CIFAR-10
ViT-Base 58.07
+ $G_{3}B$ 24 80.54
+ $G_{3}U$ 24 80.81
+ $G_{5}B$ 24 77.99
+ $G_{5}U$ 24 80.52
+ $G_{3}B$ 12 80.09
+ $G_{3}U$ 12 81.00
+ $G_{5}B$ 12 79.80
+ $G_{5}U$ 12 81.17
+ $G_{3}B$ 8 80.76
+ $G_{3}U$ 8 80.40
+ $G_{5}B$ 8 79.79
+ $G_{5}U$ 8 80.83
Table 5: Comparing performance of Fourier KArAt for ViT-Tiny and ViT-Base with or without using $\ell_{1}$ projection in Algorithm 1. ViT-Tiny and ViT-Base use softmax and do not require $\ell_{1}$ projection.
Model $\ell_{1}$ Projection Acc.@1
on CIFAR-10
ViT- Tiny 72.76
+ $G_{3}B$ 41.99
+ $G_{3}U$ 40.85
+ $G_{3}B$ 76.69
+ $G_{3}U$ 75.56
ViT-Base 83.45
+ $G_{3}B$ 47.44
+ $G_{3}U$ 46.11
+ $G_{3}B$ 80.09
+ $G_{3}U$ 81.00
B.4 Ablation Study on the Impact of grid size, $G$
KANs are highly dependent on certain hyperparameters, and Fourier KAN and Fourier-KArAt have only one hyperparameter to tune the performance — grid size $G$. Thus, we perform extensive experimentation involving grid size $G$ and present in Table 5. We observe that each of the particular ViT models, in conjunction with particular Fourier-KArAt variants, has a typical $G$ value that brings out its best performance, and there is no universal value of $G$ to follow. When performing validation with ViT-Base+ variants on CIFAR-10 and CIFAR-100, the accuracy drops as the grid size passes a size after 5. However, this behavior is not persistent with the ViT-Small/Tiny+ variants; see Table 5111We did not perform gridsize 5, 7, 9, and 11 experiments with ViT-Tiny+KArAt as it was already outperforming the base ViT-Tiny with gridsize 1 and 3, and the experiments are extensively resource intensive..
Table 6: Ablation study on the operator variants in Fourier KArAt. Note that $\widehat{\Phi}(W{\cal A})$, $W\widehat{\Phi}({\cal A})$ refer to order of using $W$ and $\widehat{\Phi}$, and $\Phi$ refers to the full-rank operator.
Model Acc.@1 on CIFAR-10
$\widehat{\Phi}(W{\cal A})$ $W\widehat{\Phi}({\cal A})$ $\Phi({\cal A})$
ViT-Base
+ $G_{11}B$ 41.02 50.47 30.14
+ $G_{11}U$ 47.74
ViT-Tiny
+ $G_{11}B$ 67.04 75.41 33.09
+ $G_{11}U$ 70.98 73.67 36.98
Model Low Rank Configurations No. of Distinct Heads Acc.@1 on
(for Fourier KArAt) softmax Fourier KArAt CIFAR-10
ViT-Base 12 0 83.45
ViT-Base + $G_{11}B$ $\hat{\sigma}=\widehat{\Phi}(W{\cal A})$ 0 12 41.02
$\hat{\sigma}=\widehat{\Phi}(W{\cal A})$ 6 6 47.60
$\hat{\sigma}=W\widehat{\Phi}({\cal A})$ 0 12 50.47
$\hat{\sigma}={\Phi}({\cal A})$ 0 12 30.14
$\hat{\sigma}={\Phi}({\cal A})$ 6 6 30.24
ViT-Tiny 3 0 72.76
ViT-Tiny + $G_{11}B$ $\hat{\sigma}={\Phi}({\cal A})$ 0 3 33.09
$\hat{\sigma}={\Phi}({\cal A})$ 1 2 32.51
$\hat{\sigma}={\Phi}({\cal A})$ 2 1 34.73
ViT-Tiny + $G_{11}U$ $\hat{\sigma}={\Phi}({\cal A})$ 0 3 36.98
$\hat{\sigma}={\Phi}({\cal A})$ 1 2 33.12
$\hat{\sigma}={\Phi}({\cal A})$ 2 1 70.57

### v2

@@ Appendix B Addendum to Empirical Study
Table 5: Hyper-parameter settings for image classification and recognition experiments conducted in this work for vanilla ViTs and ViT+Fourier KArAts.
@@
Patch Size 16
(a) Attention matrix ${\mathcal{A}^{i,j}}$ before softmax activation.
Figure 6: Spectral analysis of the attention matrices before and after softmax shows that they are low-rank. For this experiment, we use all 3 heads in the last encoder block of ViT-Tiny on 5 randomly sampled images from the CIFAR-10 validation set. We plot all 15 singular vectors (each of 197 dimensions) where the singular values are arranged in non-increasing order.
B.1 Implementation Details
We implement the framework in Python using PyTorch [62] and use the training strategy of DEiT [73]. We train all models for $100$ epochs with the ADAM optimizer [37] with a base learning rate of $3.125\times 10^{-5}$ and batch-size $32$, except for the experiments on ImageNet-1K that use a learning rate of $1.25\times 10^{-4}$ with batch-size $128$. All experiments use a warm-up of $5$ epochs and a cosine scheduler with a weight decay of $5\times 10^{-2}$. The experiments were performed on two 80 GB NVIDIA H100 GPUs. The hyperparameter settings for experiments with ViTs are given in Table 5.
B.1.1 Baselines and Datasets
For our image classification benchmarking, we chose 5 popular vision Transformers—ViT-Tiny, ViT-Small [76, 68], ViT-Base [24], ConViT-Tiny [20], and Swin-Transformer-Tiny [49]. We incorporate Fourier-KArAt in them by replacing their softmax function with a learnable activation. We perform our benchmarking on CIFAR-10, CIFAR-100 [40] and ImageNet-1K [21] datasets. We use small-scale datasets, SVHN [54], Oxford Flowers-102 [59], and STL-10 [18] for understanding the transfer learning capability of the models. Additionally, we use ViT-Det [46] on the MS COCO [48] dataset for the object detection and segmentation task.
B.1.2 Small-Scale Datasets used
We use 3 small-scale datasets, SVHN [54], Oxford Flowers 102 [59], and STL-10 [18] for transfer learning task. SVHN is similar to MNIST [43], a 10-class digit classification dataset with small images like CIFAR-10 [40], divided into 73,257 training and 26,032 test images. The challenge in this dataset comes from the distracting elements surrounding the concerned class. The OxfordFlowers-102 [59] is a small yet challenging dataset with unbalanced training images (40-254 images) for the 102 classes. The STL-10 [18] is a CIFAR-like 10-class dataset with a larger image size, but only 500 training and 800 testing images per class.
B.1.3 Implementing Different Basis and Base Activation Functions in KArAt
[…]
$\textstyle{\widehat{\phi}^{i,j}_{pq}({\cal A}^{i,j}_{k,q})=b({\cal A}^{i,j}_{k,q})+\sum_{m=1}^{G}a_{pqm}\cos{(m{\cal A}^{i,j}_{k,q})}+b_{pqm}\sin{(m{\cal A}^{i,j}_{k,q})}}.$ (6)
For the experiments in the main paper, in the general form, (6), we use $b(x)=0$ for all $x\in\mathbb{R}$, grid size, $G=3$, and the row dimension of the ${\widehat{\Phi}^{i,j}}$ is set to $r=12$, or otherwise noted. We encourage the reader to explore other configurations that differ from those previously outlined.
(i) Different Basis Functions Used and Their Results. We incorporate a variety of widely used KAN activation functions into KArAt. By design, KArAt can utilize any basis function for activating the attention units. Therefore, in addition to the Fourier basis, we embed 5 different wavelet bases, and the Rational Function basis into KArAt. The function representations for these choices of bases are outlined in Table 6.
We use the safe Padé approximation unit (PAU) of order $(m,n)$ for rational function basis; KAT also uses the same basis function [80]. In contrast to PAUs, safe PAUs ensure stability and the prevention of poles, making it a clear choice for rational functions. We only consider safe PAUs of order $(5,4)$ in this paper; other order combinations can be used.
Being an integral part in signal analyses and image reconstruction, wavelet functions serve as a candid choice for approximating signals and feature extraction [10]. Having zero mean and finite energy drives wavelets to make fair and meaningful representations of signals in local regimes. Similar to the Fourier transform, the inverse continuous wavelet transform (CWT) resolves the original signal. We consider mother wavelets, mentioned in Wav-KAN [10], as basis functions. Unique to this choice, KArAt learns scale, $s$, translation, $\tau$, as measured in the CWT, and a learnable weight coefficient $w$. For any choice of wavelet base function, $\phi$, the input is first shifted by the scale and translation parameters and then activated by the mother wavelet.
Embedding different bases in other ViT variants with KArAt is straightforward. Although our original design in Algorithm 2 represents each matrix entry $\widehat{\phi}_{pq}$ using the Fourier basis, any functions outlined in Table 6 can be used. All other components in Algorithm 2 stay the same.
We implement rational and wavelet basis activation functions from Table 6 in ViT-Tiny+KArAt and ViT-Base+KARAT using the blockwise configuration and show their performance on the image classification task on CIFAR-10 and CIFAR-100; see the results in Table 8. Performance across model variants is not consistent with the choice of wavelet basis function used. ViT-Tiny+KArAt achieves its highest Top-1 accuracy with DOG (derivative of Gaussian) on CIFAR-10 and -100, whereas Rational (5,4) breaks this trend with ViT-Base+KArAt. Another observation shows that ViT-Base+KArAt performs poorly with both Meyer and Morlet functions, contrary to ViT-Tiny+KArAt. Hence, we cannot make a proper remark regarding the choice of basis.
(ii) Base Activation Functions Used and Their Results. We implement different base functions in ViT-Tiny+KArAt using the blockwise and universal configurations and show their performance on the image classification task on CIFAR-10; see the results in Table 8. For the majority of our experiments, including those in the main paper, we define and use ZeroModule and set the base activation function to be identically zero, or $b(x)=0$. Slightly lower computations within a single KAN layer help motivate us to use this base activation function. Overall, choosing between the activations SiLU and GELU for ViT-Tiny+KArAt with $G_{3}B,G_{3}U$ modes results in a slightly lower Top-1 accuracy compared to the results using the ZeroModule in the main paper (see Table 2).
Table 6: Different basis functions and their representations for $b(x)=0$. * For our experiments, the morlet central frequency hyperparameter is $\omega_{0}=5$, but other nonnegative values can be used. ** Meyer is defined to be $(m\circ\nu)(\tilde{x})=m(\nu(\tilde{x}))$, where $m(t)=\mathbb{I}\left(-\infty,\frac{1}{2}\right](t)+\mathbb{I}{\left(\frac{1}{2},1\right)}(t)\cos\left(\frac{\pi}{2}\nu(2t-1)\right)$ and $\nu(x)=x^{4}(35-84x+70x^{2}-20x^{3})\mathbb{I}{[0,1]}(x)$.
Basis Function Representation, $\phi(x)$ Initialized Specifications
Fourier $\sum\limits_{k=1}^{G}\big{(}a_{k}\cos{(kx)}+b_{k}\sin{(kx)}\big{)}$ $a_{k},\ b_{k}\sim\mathcal{N}(0,1),\ G\text{ denotes the grid size}$
Rational $(m,n)$ $\frac{a_{0}+a_{1}x+\cdots+a_{m}x^{m}}{1+|b_{1}x+\cdots+b_{n}x^{n}|}$ $a_{i},\ b_{j}\sim\mathcal{N}(0,1),\ i=0,\ldots,m,\ j=1,\ldots,n$
Mexican Hat $\frac{2w}{\pi^{1/4}\sqrt{3}}\left(\tilde{x}^{2}-1\right)e^{-\frac{\tilde{x}^{2}}{2}},\ \tilde{x}=\frac{x-\tau}{s}$ $w\sim\mathcal{N}(0,1),\ \text{(Translation)~{}}\tau=0,\ \text{(Scale)~{}}s=1$
Morlet* $w\cos{\left(\omega_{0}\tilde{x}\right)}e^{-\frac{\tilde{x}^{2}}{2}},\ \tilde{x}=\frac{x-\tau}{s}$
DOG $-w\frac{d}{dx}\left(e^{-\frac{\tilde{x}^{2}}{2}}\right),\ \tilde{x}=\frac{x-\tau}{s}$
Meyer** $w\sin{\left(\pi|\tilde{x}|\right)}\left(m\circ\nu\right)(\tilde{x}),\ \tilde{x}=\frac{x-\tau}{s}$
Shannon $w\text{sinc}\left(\frac{\tilde{x}}{\pi}\right)\omega(\tilde{x}),\ \tilde{x}=\frac{x-\tau}{s}$ $\omega(\tilde{x})\ \text{is the symmetric hamming window}$
Table 7: Blockwise and universal ViT-Tiny+KArAt and their Top-1 accuracies on CIFAR-10 using different base activation functions (nn.SiLU, nn.Identity, nn.GELU).
Model Base Activation Acc.@1
on CIFAR-10
ViT-Tiny+$G_{3}B$ Identity 75.90
SiLU 76.25
GELU 76.48
ViT-Tiny+$G_{3}U$ Identity 75.21
SiLU 75.25
GELU 75.28
ViT-Tiny+$G_{1}B$ Identity 74.87
SiLU 75.53
GELU 75.18
ViT-Tiny+$G_{1}U$ Identity 74.44
SiLU 75.12
GELU 74.82
Table 8: Blockwise configuration ViT+KArAt and their Top-1 accuracies using different base functions on CIFAR-10 and CIFAR-100.
Model Basis Function Acc.@1
CIFAR-10 CIFAR-100
ViT-Tiny+KArAt Rational (5,4) 69.22 37.93
Mexican Hat 69.62 38.92
Morlet 70.84 37.88
DOG 73.89 44.23
Meyer 71.13 41.65
Shannon 71.41 39.40
ViT-Base+KArAt Rational (5,4) 72.59 41.52
Mexican Hat 72.58 44.44
Morlet 24.19 6.22
DOG 72.28 48.69
Meyer 28.83 6.67
Shannon 69.84 39.98
B.1.4 Deploying Fourier KArAt to other ViTs
We consider three modern ViT-based architectures, including ConViT [20] and Swin Transformer [49]. Incorporating the Fourier KArAt variants into these models is not straightforward. Below, we outline their implementation details.
(i) ConViT [20] uses two different attention mechanisms, the gated positional self-attention (GPSA) and MHSA. All of the variants of ConViT vary the number of attention heads per encoder while keeping the number of encoders the same. The purpose of the ConViT is to combine a CNN and a ViT in one encoder block. With the first ten layers using GPSA and the last two layers using MHSA, the model learns a balanced structure between CNNs and ViTs. Our variant of ConViT replaces the MHSAs with Fourier KArAts. We implement the universal and blockwise attention modes on this model. We give the ConViT’s hyperparameter configuration details in Table 9.
(ii) Swin Transformers [49] has a variable number of tokens. The number of tokens considered in the attention operation varies based on the window size, and thus, we do not implement a universal attention mode due to incompatibility. We replace regular and shifted window MHSAs, W-MHSA and SW-MHSA, with regular and shifted Fourier KArAts. We give the Swin-Transformer’s hyperparameter configuration details in Table 10.
B.1.5 Implementing Fourier KArAt for Object Detection and Semantic Segmentation
Implementing Fourier KArAt for detection and segmentation is an involved task. Firstly, RCNN [33, 30, 64] frameworks employ an augmentation strategy to train on dynamic image sizes, which also dynamically change the number of tokens and thus, token interactions in the attention. However, KArAt needs to ensure fixed-length input and output; hence, the dynamic resize becomes incompatible. Secondly, vanilla ViT-Det [46] involves window partitioning in the self-attention, making the number of tokens variable in the attention matrices throughout the blocks of encoders. Thirdly, as this window mechanism changes the number of token interactions in the attention matrices, the universal mode of KArAt also becomes incompatible. To solve all these, we restrict the input image size to $224\times 224$, and use a window size of $14$ and a patch size of $16$. This inherently fixes the number of tokens to $196$, and the window size of $14$ ensures $196$ tokens in the windowed attention blocks. Thus, the number of tokens throughout the model remains fixed. We note that the restriction on augmentation and limiting the input size impacts the baseline performance of ViT-Det. The performance analysis is detailed in §5.1 and inference visualization is provided in Figure 9.
Table 9: Hyper-parameter settings for image classification conducted in this work for vanilla ConViT-Tiny and ConViT-Tiny+Fourier KArAt.
      Input Size       $224\times 224$
      Batch Size       128 for Imagenet-1K & 32 for CIFAR-10
      Optimizer       AdamW
      Optimizer Epsilon       $1\times 10^{-8}$
      Momentum       $0.9$
      Weight Decay       $0.05$
      Gradient Clip       $1.0$
      Learning Rate Schedule       Cosine
      Learning Rate       $5\times 10^{-4}\times\frac{\text{Batch Size}}{512}$
      Warmup LR       $1\times 10^{-6}$
      Min LR       $1\times 10^{-5}$
      Epochs       100
      Decay Epochs       30
      Warmup Epochs       5
      Decay Rate       $0.1$
      Exponential Moving Average (EMA)       False
      EMA Decay       $0.99996$
      Locality up to Layer       10
      Locality Strength       1
      Random Resize & Crop Scale and Ratio       $(0.08,1.0),(0.67,1.5)$
      Random Flip       Horizontal $0.5$; Vertical $0.0$
      Color Jittering       $0.4$
      Auto-augmentation       rand-m9-mstd0.5-inc1
      Mixup       True
      Cutmix       True
      Mixup, Cutmix Probability       $0.5$, $0.5$
      Mixup Mode       Batch
      Label Smoothing       $0.1$
Table 10: Hyper-parameter settings for image classification conducted in this work for vanilla Swin-Transformer-Tiny and Swin-Transformer-Tiny+Fourier KArAt.
      Input Size       $224\times 224$
      Crop Ratio       $0.9$
      Batch Size       128 for Imagenet-1K & 32 for CIFAR-10
      Optimizer       AdamW
      Optimizer Epsilon       $1\times 10^{-8}$
      Momentum       $0.9$
      Weight Decay       $0.05$
      Gradient Clip       $1.0$
      Learning Rate Schedule       Cosine
      Learning Rate       $5\times 10^{-4}\times\frac{\text{Batch Size}}{512}$
      Warmup LR       $1\times 10^{-6}$
      Min LR       $1\times 10^{-5}$
      Epochs       100
      Decay Epochs       30
      Patience Epochs       10
      Warmup Epochs       5
      Decay Rate       $0.1$
      Exponential Moving Average (EMA)       False
      EMA Decay       $0.99996$
      Fused Layer Norm       False
      Fused Window Process       False
      Window Size       7
      Random Resize & Crop Scale and Ratio       $(0.08,1.0),(0.67,1.5)$
      Random Flip       Horizontal $0.5$; Vertical $0.0$
      Color Jittering       $0.3$
      Auto-augmentation       rand-m9-mstd0.5-inc1
      Mixup       True
      Cutmix       True
      Mixup, Cutmix Probability       $0.5$, $0.5$
      Mixup Mode       Batch
      Label Smoothing       $0.1$
(a) ViT-Tiny
(b) ViT-Small
(c) ViT-Base
Figure 7: Training loss and test accuracy of vanilla ViTs and their Fourier KArAt versions on CIFAR-10, CIFAR-100, and ImageNet-1K datasets (left to right).
Table 11: Hyper-parameter settings for object detection and instance segmentation experiments.
Input Size $224\times 224$
Batch Size 32
Optimizer AdamW
Optimizer Epsilon $1\times 10^{-6}$
Momentum $0.9,0.999$
Weight Decay $0.1$
Learning Rate Schedule Warmup Scheduler
Learning Rate $1\times 10^{-4}$
Iterations 184375
Decay Rate $0.988$
Random Flip Horizontal $0.5$; Vertical $0.0$
Patch Size 16
Attention Window Size 14
Window Attention Block Indices $0,1,3,4,6,7,9,10$
Encoder Output Layer Index $11$
Pyramid Scale Factors $4.0,2.0,1.0,0.5$
Output Channels 256
Proposal Generator Input Layers (corresponding to feature pyramid) $p2,p3,p4,p5,p6$
Proposal Generator Input Sizes $32,64,128,256,512$
Proposal Generator Training Pre-NMS Top-K $12000$
Proposal Generator Evaluation Pre-NMS Top-K $6000$
Proposal Generator Training Post-NMS Top-K $2000$
Proposal Generator Evaluation Post-NMS Top-K $1000$
Proposal Generator NMS Threshold $0.7$
Region-of-Interest Heads IOU Threshold $0.5$
Region-of-Interest Score Threshold $0.05$
Region-of-Interest NMS Threshold $0.5$
B.1.6 Plotting Loss Landscapes
Following [45], we perform principal component analysis (PCA) of the parameter change over training progression to understand the major directions of parameter convergence. Considering the fully trained model as the minimum in the loss hyperplane and the two principal component directions as $X$ and $Y$ axes, we plot the loss values over the validation set of CIFAR-10 along the $Z$ axis for ViT-Tiny and -Base for the traditional MHSA and KArAt ($G_{3}B$ and $G_{1}B$, respectively).
B.2 Computation Time, FLOPS, Memory Requirement, and Throughput of Fourier KArAt
@@
The overall computation for Fourier KArAt variants is higher than their conventional softmax MHSA, and we have delineated it in Figure 8. Primarily, the Fourier KArAt variants have a longer training time. We show the training time comparison between the traditional MHSA and its Fourier KArAt versions for 100 epochs on CIFAR-10, CIFAR-100, and ImageNet-1K datasets for all the models (ViT-Tiny, -Small, and -Base) in Figures 8– 8, respectively. We only compare the best-performing Fourier KArAt models ($G_{1}B$ and $G_{1}U$ for ViT-Base, and $G_{3}B$ and $G_{3}U$ for ViT-Tiny & ViT-Small) with their traditional softmax MHSA counterparts. We also observe that universal mode $G_{n}U$ training times are consistently slightly less than the blockwise modes $G_{n}B$.
@@
During the training, we monitored the GPU memory requirements, and as expected, Fourier KArAt variants utilize significantly more memory than traditional MHSA. In particular, the GPU memory requirements scale by 2.53×2.5-3$\times$2.5 - 3 ×, compared to the traditional softmax MHSA.
@@
We also compare the throughput during inference in Figure 8 and see slightly faster inference in universal mode than blockwise, except for ViT-Base. While there is a massive training time discrepancy between vanilla ViTs and Fourier KArAt ViTs, the inference speeds for Fourier KArAt variants are comparable to their vanilla counterparts. Although there is a minor difference in throughput between universal and blockwise modes during inference, theoretically, both variants for any model with the same grid size should have the same number of FLOPs.
@@
Figure 8: A detailed comparison of computing requirements.We compare the training times for 100 epochs with the hyperparameter settings given in Table 5 for all the datasets CIFAR-10, CIFAR-100, and ImageNet-1K. We also compare the throughputs of different models on ImageNet-1K; the throughput results will be similar for other datasets as the input size is $224\times 224$.
Table 12: Parameter, computation, and memory requirement for Fourier-KArAt (with hidden dimension, $r=12$) compared to the traditional softmax attention. This Table particularly shows the individual computation required for the attention activation. The memory requirement shown is approximate and is based on averages of batches of $32$ images of resolution $224\times 224$. We note that changing $r$ will affect the performance and memory requirements. In our main paper, all the experiments were performed with $r=12$.
Model Parameters GFLOPs GPU Memory
Attention Activation Total Attention Activation Total
ViT-Base 0 85.81M 0.016 17.595 7.44 GB
+ $G_{1}B$ 1.70M 87.51M 0.268 17.847 17.36 GB
+ $G_{1}U$ 0.14M 85.95M 0.268 17.847 16.97 GB
ViT-Small 0 22.05M 0.008 4.614 4.15 GB
+ $G_{3}B$ 1.53M 23.58M 0.335 4.941 11.73 GB
+ $G_{3}U$ 0.13M 22.18M 0.335 4.941 11.21 GB
ViT-Tiny 0 5.53M 0.005 1.262 2.94 GB
+ $G_{3}B$ 0.76M 6.29M 0.168 1.425 7.48 GB
+ $G_{3}U$ 0.06M 5.59M 0.168 1.425 7.29 GB
(a) ViT-Det with softmax attention
(b) ViT-Det Fourier KArAt $G_{1}B$
(c) ViT-Det Fourier KArAt $G_{1}U$
Figure 9: Detection and segmentation tasks inference visualization using ViT-Det with ViT-Base as backbone for traditional MHSA and Fourier KArAt. For each sample, the ground-truth is given on the right side and the inference is on the left.
(a) Performance on Object Detection
(b) Performance on Instance Segmentation
Figure 10: Training Curves of ViT-Det and Fourier KArAt in object detection and instance segmentation tasks while training from ImageNet-1K pretrained ViT-Base with traditional MHSA weights initialization. In this particular task, the universal configuration ($G_{1}U$), performs strictly better than the blockwise configuration ($G_{1}B$).
B.3 Ablation Study with KAN and Fourier KArAt
In this section, we perform a detailed ablation study with different hyperparameter settings. Our first set of experiments shows why B-Splines are not a good basis for KAN for image classification tasks. The experiments with B-Splines solidify our argument for why we primarily use the Fourier basis in KArAt. After that, we perform rigorous ablation studies of Fourier KArAt on the hidden dimension, grid size effects, and many others.
B.3.1 Experiments with KANs Operated on B-Spline Basis
We test the performance of the B-spline KANs in the classification task on the CIFAR-10 [40], CIFAR-100 [40], MNIST [42, 43], and Fashion-MNIST [77] datasets. We performed extensive experiments to see if there are any benefits to choosing B-Splines for the basis functions in KAN layers. Table 13 shows the Top-1 accuracies from different variants of a Deep KAN. In this set of experiments, we closely followed earlier works [81, 51]. Hyperparameters involved in these experiments include the order of the B-spline $k$, grid size $G$, grid range $[-I,I]$, layers, and width. Although the KANs with B-Spline basis yield high accuracies in the small-scale MNIST and Fashion-MNIST datasets, they fail to generalize over larger datasets (CIFAR-10 and -100).